How do I find increses of numbers by percents? For example; Increase 120 by 70%? How would I do that? Thank you!
Answer by apachemexican1(64)

set up the problem
x/120=70/100
now multiply 120 by 70
120*70=8400
now divide 8400 by 100
8400/100=84
the increase is 84
to find the total of the increase
just add 120 and 84
120+84=204
good luck!